# Relaxation and stability analysis of a third-order multiclass traffic flow model

### v1

@@ 1 Introduction
A key aspect in the development and understanding of traffic models is the ability to link different scales through suitable scaling limits, allowing for a consistent transition from microscopic dynamics to macroscopic behavior [27, 7, 9, 16, 10]. In this work, we focus on the microscopic and macroscopic scales, which represent the finest and coarsest levels of description, respectively.
@@
The hysteresis model we consider is inspired by the formulation introduced by Corli and Fan [13, 14], which captures the memory-dependent nature of driver behavior, particularly in congested or stop-and-go traffic regimes. We argue that modeling hesitation through hysteresis provides a more accurate representation of real traffic dynamics. In particular, it accounts for the asymmetry in driver reactions. Empirical observations suggest that drivers tend to accelerate more cautiously than they decelerate, even under identical traffic conditions – a behavior that cannot be reproduced by hesitation laws depending solely on the current density. Then, we will introduce relaxation terms in the third-order model as in the inhomogeneous ARZ model, where a relaxation term allows drivers to achieve the equilibrium speed [24]. In the small relaxation limit, the ARZ model approaches the LWR model, which can be obtained by means of a Chapman-Enskog-type expansion. Here, the stability and wellposedness of solutions to the inhomogeneous ARZ model are governed by the study of the sign of the diffusion coefficient. More precisely, the sub-characteristic condition [30] holds if and only if the advection-diffusion equation is dissipative, which is obtained via a first-order Chapman-Enskog-type expansion, provided that the system is of size two [12, Th. 3.1]. This statement, however, does not necessarily hold for hyperbolic relaxation problems of general size.
@@
Indeed, relaxation and stability results are well-established for second-order systems, whereas results for higher-order models are still partial. In fact, the local equilibrium approximation of a system of size three may not even be hyperbolic [12, Sec. 2]. Whenever a loss of hyperbolicity does not arise in the local equilibrium approximation, it makes sense to seek a first-order correction. However, it is generally an open question whether the first-order correction is dissipative even if the characteristic speeds of the local equilibrium are not excessive [12].
@@
We contribute to this research gap by conducting a stability analysis for various relaxation terms. In particular, we examine the relaxation to the ARZ model in terms of a conservative and non-conservative form. We obtain, for the conservative form of our proposed model of size three, stabilization results that are similar to those of [12, Th. 3.1], which hold only for second-order systems. Namely, the first-order correction leads to a dissipative advection-diffusion equation if and only if the characteristic curves of the local equilibrium do not exceed those of the relaxed system. On the other hand, a relaxation with respect to a non-conservative form results in a system that is of Hamilton-Jacobi-type. Although a diffusion coefficient cannot be investigated in this case, we still obtain results in terms of the characteristic speeds. A similar problem arises when the system is relaxed to a family of velocity curves. Then, a Chapman-Enskog-type expansion leads to a non-symmetric diffusion matrix, where the equivalence [12, Th. 3.1] also does not hold. In this case, we obtain a relation between the relaxed system and the zero-relaxation limit in the sense of vanishing second-order terms.
@@ 2.1 Follow-the-Leader models
[…]
@@
$-\dot{v}_{i}=-c_{\gamma}\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{\gamma+1}}=c_{\gamma}\frac{1}{(x_{i+1}-x_{i})^{\gamma-1}}\bigg{(}-\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{2}}\bigg{)}=P^{\prime}(\rho_{i})\dot{\rho}_{i}=\dot{P}(\rho_{i}),$
which implies $\dot{\omega}_{i}=0$ for $\omega_{i}\coloneqq v_{i}+P(\rho_{i})$. Since the quantity $\omega_{i}(t)=\omega_{i}(0)$ is constant over time, it is a driver-dependent quantity that corresponds to car $i$.
@@ 2.2 Multiclass models
$\dot{x}_{i}&=v_{i},\\ \dot{\omega}_{i}&=0,\\ \dot{h}_{i}&=0\end{aligned}\end{cases}\text{with}\quad\ \omega_{i}\coloneqq v_{i}+q(\rho_{i},h_{i}).$ (2.2)
@@ 2.3 The third-order model (TOM)
$\dot{x}_{i}&=v_{i},\\ \dot{\omega}_{i}&=0,\\ \dot{h}_{i}&=\dot{\textup{{h}}}(\rho_{i})\end{aligned}\end{cases}\text{with}\qquad\omega_{i}\coloneqq v_{i}+q(h_{i}).$ (2.3)
@@
we observe that there are no restrictive assumptions on the functions $q$ and h required, which must be specified for the proposed ansatz in (2.3), with the exception of the non-restrictive assumption ${q^{\prime}(h_{i})\textup{{h}}^{\prime}(\rho_{i})>0}$. Therefore, the requirement of a closure does not occur for the choice $\omega_{i}=v_{i}+q(h_{i})$, which we propose. However, the resulting ODE $\big{\{}\dot{x}_{i}=v_{i},\,\dot{\omega}_{i}=0\big{\}}$ must be closed by an additional equation, namely $\dot{h}_{i}=\dot{\textup{{h}}}(\rho_{i})$. For this reason, the model (2.3) arises rather from a modeling step of the dynamics than from a closure mapping or parameterization as in previous models. In particular, a reasonable choice for $\textup{{h}}(\rho)$ is given in [13, 14], where $h$ accounts for the hesitation of drivers.
@@ 3 The macroscopic TOM model
$\partial_{T}\tau(T,X)-\partial_{X}v(T,X)\ \ &=0,\\ \partial_{T}\Big{(}v(T,X)+q\big{(}h(T,X)\big{)}\Big{)}&=0,\\ \partial_{T}\Big{(}h(T,X)-\textup{{h}}\big{(}\rho(T,X)\big{)}\Big{)}&=0\end{aligned}\end{cases}\text{for}\quad\tau\coloneqq\frac{1}{\rho}.$ (3.4)
@@
Note that the variable $X$ is not a mass, but describes the total length occupied by cars up to the point in space $x$, since the local density $\rho$ is a dimensionless quantity.
@@
$&\text{with}&v\big{(}\tau(t,X);X\big{)}&\coloneqq v_{0}(X)+q\big{(}h_{0}(X)\big{)}-q\Big{(}h_{0}(X)-\textup{{h}}\big{(}\rho_{0}(X)\big{)}+\textup{{h}}\big{(}\tau(t,X)^{-1}\big{)}\Big{)}\\ &\text{and}&v^{\prime}\big{(}\tau(t,X);X\big{)}&\;=q^{\prime}\Big{(}h_{0}(X)-\textup{{h}}\big{(}\rho_{0}(X)\big{)}+\textup{{h}}\big{(}\tau(t,X)^{-1}\big{)}\Big{)}\textup{{h}}^{\prime}\big{(}\tau(t,X)\big{)}\tau(t,X)^{-2}.\end{aligned}$
@@
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big{(}v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big{)}\partial_{x}v&=0,\\ \partial_{t}h+v\partial_{x}h&=\partial_{t}\textup{{h}}(\rho)+v\partial_{x}\textup{{h}}(\rho).\end{aligned}\end{cases}$ ($\mathcal{S}1$)
@@
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}\Big{(}v+q(h)\Big{)}+v\partial_{x}\Big{(}v+q(h)\Big{)}&=0,\\ \partial_{t}\Big{(}h-\textup{{h}}(\rho)\Big{)}+v\partial_{x}\Big{(}h-\textup{{h}}(\rho)\Big{)}&=0\end{aligned}\end{cases}$ ($\mathcal{S}2$)
@@
[…]
@@
00 0 $=\Big{(}\partial_{t}\rho+\partial_{x}(\rho v)\Big{)}\Big{(}h-\textup{{h}}(\rho)\Big{)}+\rho\partial_{t}\Big{(}h-\textup{{h}}(\rho)\Big{)}+\rho v\partial_{x}\Big{(}h-\textup{{h}}(\rho)\Big{)}$
@@
00 0 $=\partial_{t}\Big{(}v+q(h)\Big{)}+v\partial_{x}\Big{(}v+q(h)\Big{)}+\Big{(}v+q(h)\Big{)}\Big{(}\partial_{t}\rho+\partial_{x}(\rho v)\Big{)}$
@@
Although two eigenvalues coincide, the Jacobian of the systems ($\mathcal{S}1$) – ($\mathcal{S}3$) have a full set of eigenvectors and hence form a strongly hyperbolic system. In fact, systems that are endowed with an entropy-entropy flux pair can be symmetrized and hence are strongly hyperbolic [22, Th. 3.1].
Furthermore, we recall that a characteristic field, corresponding to an eigenvalue $\lambda_{k}$, is genuinely nonlinear if the function $\lambda_{k}(\textup{{u}})\cdot r_{k}(\textup{{u}})$ u ) never vanishes and it is linearly degenerate if the equation $\nabla_{\textup{{u}}}\lambda_{k}(\textup{{u}})\cdot r_{k}(\textup{{u}})=0$ holds for all states u. We observe the expressions $\nabla_{\textup{{u}}}\lambda_{2}(\textup{{u}})\cdot r_{2}(\textup{{u}})=0$, $\nabla_{\textup{{u}}}\lambda_{3}(\textup{{u}})\cdot r_{3}(\textup{{u}})=0$ and the strict equality
@@
provided that the properties $\textup{{h}}^{\prime}(\rho)>0$, $q^{\prime}(h)>0$, $\textup{{h}}^{\prime\prime}(\rho)\geq 0$ and $q^{\prime\prime}(h)\geq 0$ hold. Hence, the fields $k=2,3$ that correspond to the double eigenvalue $\lambda_{v}(v)=v$, which travel faster then the first field, i.e. $\lambda_{1}(\rho,v,h)<\lambda_{v}(v)$, form contacts. Riemann invariants $\mathcal{R}_{k}$ satisfy $\nabla_{\textup{{u}}}\mathcal{R}_{k}(\textup{{u}})\cdot\lambda_{k}=0$ and read as $\mathcal{R}_{1}(\textup{{u}})=\omega=v+q(h)$ and $\mathcal{R}_{2}(\textup{{u}})=v$, $\mathcal{R}_{3}(\textup{{u}})=v$.
@@ 4.1 Relaxation in the Aw-Rascle-Zhang model
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}\Big{(}v+P(\rho)\Big{)}+v\partial_{x}\Big{(}v+P(\rho)\Big{)}&=\frac{1}{\varepsilon}\Big{(}V_{\textup{eq}}(\rho)-v\Big{)},\end{aligned}\end{cases}$ (4.7)
@@
$\dot{x}_{i}&=v_{i},\\ \dot{v}_{i}&=c_{\gamma}\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{\gamma+1}}+\frac{1}{\varepsilon}\Big{(}V_{\text{eq}}(\rho_{i})-v_{i}\Big{)},\end{aligned}\end{cases}$ (4.8)
@@
$\partial_{t}\rho+\partial_{x}\Big{(}z-\rho q\big{(}\textup{{H}}(\rho)\big{)}\Big{)}&=0,\\ \partial_{t}z+\partial_{x}\Big{(}\frac{z^{2}}{\rho}-zq\big{(}\textup{{H}}(\rho)\big{)}\Big{)}&=\frac{\rho}{\varepsilon}\Big{(}V_{\textup{eq}}(\rho)-v(\rho,z)\Big{)}\end{aligned}\end{cases}\text{for}\ \ \ v(\rho,z)=\frac{z}{\rho}-q\big{(}\textup{{H}}(\rho)\big{)}.$ (4.9)
@@
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big{(}v-q^{\prime}\big{(}\textup{{H}}(\rho)\big{)}\textup{{H}}^{\prime}(\rho)\rho\Big{)}\partial_{x}v&=\frac{1}{\varepsilon}\Big{(}V_{\textup{eq}}(\rho)-v\Big{)}.\end{aligned}\end{cases}$ (4.10)
@@
with positive diffusion $\text{with positive diffusion coefficient}\quad\mu(\rho)\coloneqq-\,\rho^{2}V_{\textup{eq}}^{\prime}(\rho)\big{(}V_{\textup{eq}}^{\prime}(\rho)+h^{\prime}(\rho)\big{)}\geq 0$
provided that the inequality $V_{\textup{eq}}^{\prime}(\rho)+h^{\prime}(\rho)\geq 0$ holds. We emphasize that the equivalence of a sub-characteristic condition and a dissipative advection-diffusion equation, which has been established for general systems of size two [12, Th. 3.1], does not necessarily hold for hyperbolic relaxation problems of general size.

### v2

@@ 1 Introduction
A key aspect in the development and understanding of traffic models is the ability to link different scales through suitable scaling limits, allowing for a consistent transition from microscopic dynamics to macroscopic behavior [8, 11, 18, 12]. For instance, small variability of microscopic velocities at equilibrium leads to stop-and-go waves, which are macroscopically observed in traffic flow [29, 23]. In this work, we focus on the microscopic and macroscopic scales, which represent the finest and coarsest levels of description, respectively.
@@
The hysteresis model we consider is inspired by the formulation introduced by Corli and Fan [15, 16], which captures the memory-dependent nature of driver behavior, particularly in congested or stop-and-go traffic regimes. We argue that modeling hesitation through hysteresis provides a more accurate representation of real traffic dynamics. In particular, it accounts for the asymmetry in driver reactions. Empirical observations suggest that drivers tend to accelerate more cautiously than they decelerate, even under identical traffic conditions – a behavior that cannot be reproduced by hesitation laws depending solely on the current density. Then, we will introduce relaxation terms in the third-order model as in the inhomogeneous ARZ model, where a relaxation term allows drivers to achieve the equilibrium speed [26]. In the small relaxation limit, the ARZ model approaches the LWR model, which can be obtained by means of a Chapman-Enskog-type expansion. Here, the stability and wellposedness of solutions to the inhomogeneous ARZ model are governed by the study of the sign of the diffusion coefficient. More precisely, the sub-characteristic condition [33] holds if and only if the advection-diffusion equation is dissipative, which is obtained via a first-order Chapman-Enskog-type expansion, provided that the system is of size two [14, Th. 3.1]. This statement, however, does not necessarily hold for hyperbolic relaxation problems of general size.
@@
Indeed, relaxation and stability results are well-established for second-order systems, whereas results for higher-order models are still partial. In fact, the local equilibrium approximation of a system of size three may not even be hyperbolic [14, Sec. 2]. Whenever a loss of hyperbolicity does not arise in the local equilibrium approximation, it makes sense to seek a first-order correction. However, it is generally an open question whether the first-order correction is dissipative even if the characteristic speeds of the local equilibrium are not excessive [14].
@@
We contribute to this research gap by conducting a stability analysis for various relaxation terms. In particular, we examine the relaxation to the ARZ model in terms of a conservative and non-conservative form. We obtain, for the conservative form of our proposed model of size three, stabilization results that are similar to those of [14, Th. 3.1], which hold only for second-order systems. Namely, the first-order correction leads to a dissipative advection-diffusion equation if and only if the characteristic curves of the local equilibrium do not exceed those of the relaxed system. On the other hand, a relaxation with respect to a non-conservative form results in a system that is of Hamilton-Jacobi-type. Although a diffusion coefficient cannot be investigated in this case, we still obtain results in terms of the characteristic speeds. A similar problem arises when the system is relaxed to a family of velocity curves. Then, a Chapman-Enskog-type expansion leads to a non-symmetric diffusion matrix, where the equivalence [14, Th. 3.1] also does not hold. In this case, we obtain a relation between the relaxed system and the zero-relaxation limit in the sense of vanishing second-order terms.
@@ 2.1 Follow-the-Leader models
$\begin{cases}\begin{aligned} \dot{x}_{i}&=v_{i},\\ \dot{v}_{i}&=c_{\gamma}\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{\gamma+1}},\end{aligned}\end{cases}$ (2.1)
@@
$\dot{v}_{i}=-c_{\gamma}\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{\gamma+1}}=c_{\gamma}\frac{1}{(x_{i+1}-x_{i})^{\gamma-1}}\bigg(-\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{2}}\bigg)=P^{\prime}(\rho_{i})\dot{\rho}_{i}=\dot{P(\rho_{i})},$
where $\dot{P(\rho_{i})}\coloneqq\frac{\textup{d}}{\textup{d}t}P\big(\rho_{i}(t)\big)$ denotes the composite time derivative. Hence, the relation $\dot{\omega}_{i}=0$ holds for $\omega_{i}\coloneqq v_{i}+P(\rho_{i})$. Since the quantity $\omega_{i}(t)=\omega_{i}(0)$ is constant over time, it is a driver-dependent quantity that corresponds to car $i$.
@@ 2.2 Multiclass models
$\begin{cases}\begin{aligned} \dot{x}_{i}&=v_{i},\\ \dot{\omega}_{i}&=0,\\ \dot{h}_{i}&=0\end{aligned}\end{cases}\text{with}\quad\ \omega_{i}\coloneqq v_{i}+q(\rho_{i},h_{i}).$ (2.2)
@@ 2.3 The third-order model (TOM)
$\begin{cases}\begin{aligned} \dot{x}_{i}&=v_{i},\\ \dot{\omega}_{i}&=0,\\ \dot{h_{i}}&=\dot{\textup{{h}}(\rho_{i})}\end{aligned}\end{cases}\text{with}\qquad\omega_{i}\coloneqq v_{i}+q(h_{i}).$ (2.3)
@@
we observe that there are no restrictive assumptions on the functions $q$ and h required, which must be specified for the proposed ansatz in (2.3), with the exception of the non-restrictive assumption $q^{\prime}(h_{i})\textup{{h}}^{\prime}(\rho_{i})>0$}. Namely, the term $q^{\prime}(h_{i})\textup{{h}}^{\prime}(\rho_{i})$ generalizes the derivative $P^{\prime}(\rho_{i})>0$ in the follow-the-leader models in Section 2.1. Therefore, the requirement of a closure does not occur for the choice $\omega_{i}=v_{i}+q(h_{i})$, which we propose. However, the resulting ODE $\big\{\dot{x}_{i}=v_{i},\,\dot{\omega}_{i}=0\big\}$ must be closed by an additional equation, namely $\dot{h}_{i}=\dot{\textup{{h}}}(\rho_{i})$. For this reason, the model (2.3) arises rather from a modeling step of the dynamics than from a closure mapping or parameterization as in previous models. In particular, a reasonable choice for $\textup{{h}}(\rho)$ is given in [15, 16], where $h$ accounts for the hesitation of drivers.
@@ 3 The macroscopic TOM model
$\begin{cases}\begin{aligned} \partial_{T}\tau(T,X)-\partial_{X}v(T,X)\ \ &=0,\\ \partial_{T}\Big(v(T,X)+q\big(h(T,X)\big)\Big)&=0,\\ \partial_{T}\Big(h(T,X)-\textup{{h}}\big(\rho(T,X)\big)\Big)&=0\end{aligned}\end{cases}\text{for}\quad\tau\coloneqq\frac{1}{\rho}.$ (3.4)
@@
Note that the variable $X$ is not a mass, but describes the total length occupied by cars up to the point in space $x$, since the local density $\rho$ is a dimensionless quantity. Likewise, the driver-dependent quantity $h$ is dimensionless, while the function $q$ has the physical dimension of a velocity.
@@
$\begin{aligned} &\text{with}&v\big(\tau(t,X);X\big)&\coloneqq v_{0}(X)+q\big(h_{0}(X)\big)-q\Big(h_{0}(X)-\textup{{h}}\big(\rho_{0}(X)\big)+\textup{{h}}\big(\tau(t,X)^{-1}\big)\Big)\\ &\text{and}&v^{\prime}\big(\tau(t,X);X\big)&\;=q^{\prime}\Big(h_{0}(X)-\textup{{h}}\big(\rho_{0}(X)\big)+\textup{{h}}\big(\tau(t,X)^{-1}\big)\Big)\textup{{h}}^{\prime}\big(\tau(t,X)\big)\tau(t,X)^{-2}.\end{aligned}$
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big(v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big)\partial_{x}v&=0,\\ \partial_{t}h+v\partial_{x}h&=\partial_{t}\textup{{h}}(\rho)+v\partial_{x}\textup{{h}}(\rho).\end{aligned}\end{cases}$ ($\mathcal{S}1$)
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}\Big(v+q(h)\Big)+v\partial_{x}\Big(v+q(h)\Big)&=0,\\ \partial_{t}\Big(h-\textup{{h}}(\rho)\Big)+v\partial_{x}\Big(h-\textup{{h}}(\rho)\Big)&=0\end{aligned}\end{cases}$ ($\mathcal{S}2$)
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}\Big(\rho v(\rho,z,w)\Big)&=0,\\ \partial_{t}z+\partial_{x}\Big(\frac{z^{2}}{\rho}-zq(h)\Big)&=0,\\ \partial_{t}w+\partial_{x}\Big(wv(\rho,z,w)\Big)&=0.\end{aligned}\end{cases}$ ($\mathcal{S}3$)
@@
$0$ $=\Big(\partial_{t}\rho+\partial_{x}(\rho v)\Big)\Big(h-\textup{{h}}(\rho)\Big)+\rho\partial_{t}\Big(h-\textup{{h}}(\rho)\Big)+\rho v\partial_{x}\Big(h-\textup{{h}}(\rho)\Big)$
@@
$0$ $=\partial_{t}\Big(v+q(h)\Big)+v\partial_{x}\Big(v+q(h)\Big)+\Big(v+q(h)\Big)\Big(\partial_{t}\rho+\partial_{x}(\rho v)\Big)$
@@
Although two eigenvalues coincide, the Jacobian of the systems ($\mathcal{S}1$) – ($\mathcal{S}3$) have a full set of eigenvectors and hence form a strongly hyperbolic system. In fact, systems that are endowed with an entropy-entropy flux pair can be symmetrized and hence are strongly hyperbolic [24, Th. 3.1].
Furthermore, we recall that a characteristic field, corresponding to an eigenvalue $\lambda_{k}$, is genuinely nonlinear if the function $\nabla_{\textup{{u}}}\lambda_{k}(\textup{{u}})\cdot r_{k}(\textup{{u}})$ never vanishes and it is linearly degenerate if the equation $\nabla_{\textup{{u}}}\lambda_{k}(\textup{{u}})\cdot r_{k}(\textup{{u}})=0$ holds for all states u. We observe the expressions $\nabla_{\textup{{u}}}\lambda_{2}(\textup{{u}})\cdot r_{2}(\textup{{u}})=0$, $\nabla_{\textup{{u}}}\lambda_{3}(\textup{{u}})\cdot r_{3}(\textup{{u}})=0$ and the strict equality
@@
provided that the properties
$\textup{{h}}^{\prime}(\rho)>0,\quad q^{\prime}(h)>0,\quad\textup{{h}}^{\prime\prime}(\rho)\geq 0,\quad q^{\prime\prime}(h)\geq 0$
hold, which is slightly more restrictive than the assumption $q^{\prime}(h_{i})h^{\prime}(h_{i})>0$ in Section 2.3, where the microscopic case is considered. Hence, the fields $k=2,3$ that correspond to the double eigenvalue $\lambda_{v}(v)=v$, which travel faster then the first field, i.e. $\lambda_{1}(\rho,v,h)<\lambda_{v}(v)$, form contacts. Riemann invariants $\mathcal{R}_{k}$ satisfy $\nabla_{\textup{{u}}}\mathcal{R}_{k}(\textup{{u}})\cdot\lambda_{k}=0$ and read as $\mathcal{R}_{1}(\textup{{u}})=\omega=v+q(h)$ and $\mathcal{R}_{2}(\textup{{u}})=v$, $\mathcal{R}_{3}(\textup{{u}})=v$.
@@
This means that shock waves are stable and propagate in the correct direction, namely backward relative to traffic flow, as observed in reality. Furthermore, the total entropy of the system is conserved, since the entropy inequality $\partial_{t}\eta+\partial_{x}\psi=0$ is in fact an equality. In contrast to gas dynamics where information is lost over a shock wave, which yields mathematically an entropy inequality, the total entropy in traffic flow models is expected to be constant. For instance, a strict entropy inequality would mean that information about the driver dependent quantity $h$ is lost across a shock [10, 31].
We note that oscillations are sometimes of interest, since they model stop-and-go waves. However, those arise from instabilities of small perturbations around steady states and not from a shock wave.
@@ 4.1 Relaxation in the Aw-Rascle-Zhang model
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}\Big(v+P(\rho)\Big)+v\partial_{x}\Big(v+P(\rho)\Big)&=\frac{1}{\varepsilon}\Big(V_{\textup{eq}}(\rho)-v\Big),\end{aligned}\end{cases}$ (4.7)
@@
$\begin{cases}\begin{aligned} \dot{x}_{i}&=v_{i},\\ \dot{v}_{i}&=c_{\gamma}\frac{\dot{x}_{i+1}-\dot{x}_{i}}{(x_{i+1}-x_{i})^{\gamma+1}}+\frac{1}{\varepsilon}\Big(V_{\text{eq}}(\rho_{i})-v_{i}\Big),\end{aligned}\end{cases}$ (4.8)
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}\Big(z-\rho q\big(\textup{{H}}(\rho)\big)\Big)&=0,\\ \partial_{t}z+\partial_{x}\Big(\frac{z^{2}}{\rho}-zq\big(\textup{{H}}(\rho)\big)\Big)&=\frac{\rho}{\varepsilon}\Big(V_{\textup{eq}}(\rho)-v(\rho,z)\Big)\end{aligned}\end{cases}\text{for}\ \ \ v(\rho,z)=\frac{z}{\rho}-q\big(\textup{{H}}(\rho)\big).$ (4.9)
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big(v-q^{\prime}\big(\textup{{H}}(\rho)\big)\textup{{H}}^{\prime}(\rho)\rho\Big)\partial_{x}v&=\frac{1}{\varepsilon}\Big(V_{\textup{eq}}(\rho)-v\Big).\end{aligned}\end{cases}$ (4.10)
@@
$\text{with positive diffusion coefficient}\quad\mu(\rho)\coloneqq-\,\rho^{2}V_{\textup{eq}}^{\prime}(\rho)\big(V_{\textup{eq}}^{\prime}(\rho)+h^{\prime}(\rho)\big)\geq 0$
provided that the inequality $V_{\textup{eq}}^{\prime}(\rho)+h^{\prime}(\rho)\geq 0$ holds. We emphasize that the equivalence of a sub-characteristic condition and a dissipative advection-diffusion equation, which has been established for general systems of size two [14, Th. 3.1], does not necessarily hold for hyperbolic relaxation problems of general size.